OpenStudy (anonymous):

if 7log base a 16/15+5log base a 25/24 +3log base a 81/80=8 then a is

OpenStudy (zzr0ck3r):

= log((16/15)^7)+log((25/24)^5)+log((81/80)^3)=8 a^[log((16/15)^7)+log((25/24)^5)+log((81/80)^3)]=a^8 a^log((16/15)^7) * a^log((25/24)^5) * a^log((81/80)^3=a^8 (16/15)^7 * (25/24)^5 * (81/80)^3 = a^8 a^8 = 2 \[a=\sqrt[8]{2}\]

OpenStudy (anonymous):

thankz

OpenStudy (zzr0ck3r):

np, sorry for typos. I think I fixed them all

OpenStudy (zzr0ck3r):

do you understand all the steps?

OpenStudy (zzr0ck3r):

properties used a*log(b) = log(b^a) and a^b*a^c = a^(b+c) and a^log_a(x) = x

OpenStudy (zzr0ck3r):

I think technically \[a=\pm \sqrt[8]{2}\]